OpenStudy (anonymous):

If I buy 2 apples and 6 oranges, I will spend $5.60. However, if I buy 6 apples and 2 oranges, I will spend $4.80. How much does one orange cost? A. $0.55 B. $0.70 C. $0.75 D. $0.60

hero (hero):

2x + 6y = 5.60 6x + 2y = 4.80

OpenStudy (anonymous):

I still don't understand.. :/ Math is so confusing to me!

OpenStudy (anonymous):

Use substitution or elimination. I would prefer to use elimination.

hero (hero):

Actually, you don't need that other part. But yeah, do what Fakshon said

OpenStudy (anonymous):

:/ Can you give an example?

OpenStudy (anonymous):

2x + 6y = 5.60 6x + 2y = 4.80 Turn get rid of the y's: 2x+6y = 5.60 (-3)6x+(-3)2y = (-3)4.80 2x+6y = 5.60 -18x-6y = -14.40 Then add the to together: -16x = -8.8 divide both by -16 x = 0.55 So an apple costs .55 Plug that in to the first equation: 2(.55)+6y = 5.60 1.10+6y=5.60 -1.10 -1.10 6y = 4.50 Divide both by 6: y = .75 So an orange costs .75

OpenStudy (anonymous):

Thank you!!